# Data Guard: A Fine-grained Purpose-based Access Control System for Large Data Warehouses

### v1

@@ 6.1. Methodology
  • Primitive Fields: columns of string type to simulate simple flat data structures,
  • Flat Structs: structs with varying number of primitive fields (from single-field structs to structs containing five primitive fields, e.g. STRUCT<field_1: STRING, …, field_5:STRING>),
  • Nested Structs: structs containing nested fields of varying depths. For example, a three-level deep struct produced by our tool has the following shape:
    nfield_l1: STRUCT<pfield: STRING, nfield_l2:
    STRUCT<pfield: STRING, nfield_l3: STRUCT<pfield: STRING$>$
  • Arrays: columns of array type, with varying entries per array (e.g. 10, 50, 100). We generate datasets containing one- and two-dimensional arrays, to simulate schemas that are commonly used in AI/ML training pipelines.
For the cluster benchmarks, we generate a synthetic dataset containing 10 million rows. The size of the dataset is chosen so that the task startup overheads for benchmarking queries get amortized over a sufficiently large partition of rows processed by each task. The data is written in ORC storage format with ZLIB compression, which is the storage format adopted for LinkedIn’s warehouse data. We note however that the performance numbers reported in this section should generalize to other commonly used formats like Parquet.
The primary performance metric is the per-row CPU time difference between queries using the masking operator (via the
$MASK\_FIELD\_IF()$ UDF) and those using an ”identity” operator (a no-op UDF that returns the input as-is) on the same field. This approach isolates the computational overhead of the masking operation, eliminating the influence of external factors like storage latency, network performance, and disk I/O.
Since the goal is to analyze performance trends rather than absolute values, this metric effectively illustrates how the masking operator overhead changes across different scenarios. Both UDFs are implemented using Spark’s Expression APIs to ensure consistency.
6.2. Experiment Results
The results shown in this section are from cluster benchmarks unless mentioned otherwise.
6.2.1. Impact of Column Size
@@
We measure the overhead of applying the $MASK\_FIELD\_IF()$ UDF on a column as a function of its size. In Figure 6, we show the CPU overhead of masking a struct column containing varying number of string fields.
Figure 6. Impact of field size on CPU overhead (cluster-benchmark)
From Figure 6, we see that the overhead for masking a primitive field (string type in this case) is slightly lower than that for complex fields such as structs. But increase in the column size does not have any impact on the overhead. We confirm this behavior through a CPU profile of the $MASK\_FIELD\_IF()$ UDF which shows that evaluating the input column is slightly more expensive for a struct compared to primitive types such as string, in the generic masking implementation.
6.2.2. Impact of Field Depth
We analyze the performance of the $MASK\_FIELD\_IF()$ UDF as a function of the depth of the field being masked. The synthetic data generation tool generates random records conforming to the deeply nested schema shown in Figure 7(a), where the struct at each level contains a string field and another nested struct. Figure 7(b) shows the overhead for masking the string fields at varying depths, starting with nfield_l1.pfield (represented by the x-axis label $d=1$) and moving deeper down the hierarchy (e.g., nfield_l1.nfield_l2.pfield).
(a) Deeply Nested Schema
(b) Impact of field path depth on CPU overhead (cluster-benchmark)
Figure 7. Impact of field path depth
As seen from Figure 7(b), the overhead grows linearly with increasing depth of the string fields. CPU profiling reveals that this behavior stems from how nested struct objects are represented internally in Spark. The struct objects are organized in a tree-like hierarchy, so the $MASK\_FIELD\_IF()$ UDF must traverse this tree to locate and transform the target field. Upon transformation of the target field, each parent struct in the tree needs to be reconstructed with the modified field due to immutability of objects passed to the UDF.
6.2.3. Impact of number of fields masked
In this section, we evaluate the masking overhead as a function of the number of fields that need to be masked. For simplicity, we consider a ”flat” struct consisting of five string fields (Figure 8(a)) and vary the number of members of the struct that are masked.
(a) A flat struct containing primitive fields
(b) Impact of number of masked fields on CPU overhead (cluster-benchmark)
Figure 8. Impact of number of maskings
As seen in Figure 8(b), the overhead increases linearly with the number of fields being masked. The schema considered in Figure 8(a) does however point to an optimization opportunity, where instead of invoking the masking operation separately for each member of the struct, the masking operations across members of the struct can be performed together at once. We expect this optimization to reduce the per-field overhead incurred due to schema traversal and struct object transformation and reconstruction. A special case of this optimization is implemented in the policy compiler, where the UDF performs struct-level masking when all its members need to be masked. We plan to incorporate the generalization of this optimization in our UDF implementation to further exploit the UDF batching feature of the policy compiler.
6.2.4. Impact of consent rates
@@
The % of rows from which one or more fields need to be masked depends on the fraction of data subjects that consent to allow access to those fields. We refer to this fraction as the consent rate.
Figure 9. Impact of consent rates on CPU overhead (micro-benchmark)
@@
Figure 9 contains the micro-benchmark results showing the CPU overhead for different data types as a function of the consent rate. A consent rate of 1 reduces the $MASK\_FIELD\_IF()$ UDF to a no-op action. As expected, the average CPU overhead increases as the consent rate decreases for each data type. As discussed in the previous sections, struct column masking has higher overhead than string column masking for all consent rates. Similarly, masking a field inside a struct has a higher overhead than struct column masking due to the additional overhead of schema traversal and object transformation.
6.2.5. Data Masking in Arrays
@@
We next evaluate the overhead of the masking operation on array-typed columns where each element is of type struct. Specifically, the column has the following data type: $acol:array(struct<field_{1}:string,field_{2}:string>)$. We consider masking overhead on two different field paths:
  1. (1)
    \$.$\$.acol.[item].field_{1}$: which indicates unconditional masking of $field_{1}$ from all elements in the $acol$, and
  2. (2)
    \$.$\$.acol[?(@.field_{1}==^{\prime}jackfruit^{\prime})].[item].field_{2}$: which indicates conditional masking of $field_{2}$ based on the value of $field_{1}$.
Figure 10. Impact of array sizes on CPU overhead (cluster-benchmark)
@@
Figure 10 shows that the masking overhead grows linearly with array size for each field path. For a given array size, the overhead of unconditional masking is smaller than conditional masking due to the additional cost of comparing element values. In our implementation, this comparison requires conversion between Spark’s UTF-8 strings to Java UTF-16 strings, which is the dominant contributor to the overhead.
7. Practical Considerations
@@
In this section, we discuss practical issues that needed to be considered when deploying Data Guard at-scale in LinkedIn.
1. View Schema Evolution:  In a large warehouse, tables often undergo changes to schema and new field additions (either top-level columns or nested fields) to tables are very common. As Data Guard data-masking views are schema-preserving, the schema of the view at creation time may be different from the schema of the view at the consumption time due to newly added fields. This means that newly added fields may accidentally be consumed in user queries, resulting in non-compliant accesses. To prevent this data leakage, we dynamically rewrite the data-masking view at the view consumption time via the ViewShift $getView()$ API to set all newly added field values to $NULL$.
2. Handling non-nullable fields:  The data-masking function described thus far replaces fields to be masked with $NULL$ values and thus, assumes that fields are nullable. In large warehouses, it is common to define schemas containing non-nullable fields. To ensure compliance, our approach when rolling out Data Guard at LinkedIn, has been to over-filter data and filter out the entire row from the result set when a non-nullable field needs to be masked. It is possible to extend this behavior to provide non-NULL replacement values which are defined system-wide. This would however require changes by data consumers to handle such values in their applications. An area of ongoing work is to enhance Data Guard’s APIs to provide data consumer-specified replacement values for non-nullable fields in order to avoid over-filtering data and maximize data utility.
3. Enforcement Verification:  In addition to proactive enforcement of policies at the time of data access, Data Guard also monitors access logs for all data accesses to the warehouse data. Data Guard’s monitoring sub-system relies on raw access logs from underlying storage system (i.e. HDFS in our case) along with query execution logs from data processing engines like Spark and Trino to detect any inadvertent access to warehouse data that bypasses Data Guard enforcement. Data Guard also detects accesses to stale versions of Data Guard views from the access logs. In each scenario, Data Guard notifies consumers about potential access violations and provides them with steps to remediate their accesses.
4. Result Caching:  Commercial warehouses commonly support result set caching, even though this functionality is not available currently in LinkedIn’s warehouse. One of the conditions for result caching is that queries should not include non-deterministic SQL functions such as CURRENT_TIMESTAMP(). Thus, the data-masking views described in this paper are not eligible for result set caching. However, caching results of data-masking views has limited value in our environment because the view evaluation depends on data subject consents which are frequently updated. In scenarios where such updates are infrequent, result set caching can indeed be effective. To enable reuse of cached results, alternatives such as defining data-masking views as parametrized views that use timestamp as a parameter passed to the view or defining the masking views as materialized views appear viable.
@@
5. Engine optimizations:  An important limitation of implementing the data-masking operator using UDFs is that it prevents engines from applying optimizations such as predicate pushdown or nested column pruning. Our primary focus in implementing and rolling out Data Guard at-scale inside LinkedIn has been on the APIs for enforcing policies and ensuring a friction-free developer experience. While we implemented a number of performance optimizations to ensure that the cost of adopting data-masking views remains acceptable, we recognize the opportunity for further optimizations in the future.
8. Conclusions And Future Work
@@
We have presented Data Guard - a fine-grained access control system for large data warehouses. Data Guard’s policy language supports specification of access-control rules based on semantic labels used to describe data present in the warehouse tables. Data Guard’s policy compiler translates these policies into SQL views which encode the data-masking logic. Further, Data Guard transparently routes table accesses to the appropriate data-masking views based on the accessor’s purpose, thus supporting purpose-based access control. Data Guard has been deployed at-scale in LinkedIn and is the primary policy enforcement mechanism for accesses to warehouse tables. There are a number of avenues for further optimizing the performance of Data Guard’s data-masking operator as highlighted in the paper. We also plan to enhance Data Guard to support additional capabilities to enable privacy-enhancing technologies such as anonymization and differential privacy, and encryption for data storage.
Acknowledgements.
@@
We would like to thank Kapil Surlaker and Raghu Hiremagalur for providing numerous critical inputs on the design of Data Guard. We would like to thank Yong Li, Manasa Subramanian, Divya Singh, Yanwen Lin, Justin Heaton, Vishwaa Patel, Bryan Ji, Qiang Fu, Steve Cao, Wenning Ding, and Carleen Li for their significant contributions to the development of Data Guard. We also thank Kiran Shivaram, Souvik Ghosh, Chris Harris, David Leung and their team members for being early adopters and providing valuable customer feedback that helped improve the product. Finally, we would like to thank Maneesh Varshney and Chris Harris for their helpful feedback in improving the quality of the paper.

### v2

@@ VI-A Micro-benchmark
  • Primitive Fields: columns of string type to simulate simple flat data structures.
  • Flat Structs: structs with varying number of primitive fields.
  • Nested Structs: structs with nested fields of varying depths.
  • Arrays: columns of array type, with varying array sizes. Each array element is a struct with multiple primitive fields.
For our benchmarks, we generate a synthetic dataset with 10M rows, stored in ORC format with ZLIB compression — the same format used in LinkedIn’s data warehouse. While the reported results are based on ORC, we expect similar trends for other widely used formats such as Parquet.
In each of the following experiments, we measure the overhead of data-masking views by running two types of queries on the generated dataset:
  1. 1.
    Baseline Query: Reads the original table directly, applying an ”identity” UDF to the target field(s). This UDF simply returns the input value unchanged, serving as a control to measure the cost of data access without masking.
  2. 2.
    Masking Query: Reads the same table but with data access policies applied on the targeted fields, resulting in a data-masking view with MASK_FIELD_IF() UDF calls on those field(s).
The overhead of data masking is computed as the relative difference in CPU time per row between the Masking Query and the Baseline Query. All experiments are run on a Spark cluster with 20 worker nodes, each with 16 vCPUs and 64 GB memory. Each experiment is repeated 5 times and the average CPU time is reported.
VI-A1 Impact of Data Type
@@
We measure the overhead of applying a data access policy on a column as a function of its size. In Figure 6, we show the CPU overhead of masking a flat struct column containing varying number of fields.
Figure 6: Impact of field size on CPU overhead
Figure 6 shows that masking a primitive field (such as a string) incurs slightly less overhead compared to masking complex fields like structs. Additionally, increasing the number of fields in the column does not significantly affect the overhead.
VI-A2 Impact of Field Depth
We measure the overhead of applying a data access policy on a field as a function of its depth within a nested struct.
Figure 7: Impact of field path depth on CPU overhead
Figure 7 demonstrates that the overhead increases linearly with the depth of the enforced fields. This is because struct objects are organized hierarchically, requiring the MASK_FIELD_IF() UDF to traverse the tree structure to locate and mask the target field. After masking, each parent struct along the path must be reconstructed to reflect the change, as objects passed to the UDF are immutable.
VI-A3 Impact of number of policies applied
In order to understand the performance of data-masking views as a function of number of policies, we vary the number of invocations of the MASK_FIELD_IF() UDF on a struct column, one invocation per field of the struct column. In other words, we assume each policy results in an invocation of the masking UDF.
Figure 8: Impact of number of applied policies on CPU overhead
Figure 8 shows that the CPU overhead increases linearly with the number of policies applied. This is because each invocation of the MASK_FIELD_IF() UDF introduces a fixed cost for schema traversal and object transformation. Notably, there is potential for optimization when multiple fields within the same struct are masked: batching these operations into a single UDF call can reduce redundant overhead. We plan to generalize this optimization in our UDF implementation to leverage the policy compiler’s batching capabilities.
VI-A4 Impact of consent rates
@@
In this experiment, we measure the overhead of applying data access policies as a function of the fraction of rows that require masking. This fraction depends on whether data subjects consent to allow access to those fields, which we refer to as the consent rate.
Figure 9: Impact of consent rates on CPU overhead
@@
Figure 9 contains the micro-benchmark results showing the CPU overhead for different data types as a function of the consent rate. A consent rate of 1.0 reduces the masking query to the baseline query. As expected, the average CPU overhead decreases as the consent rate increases for each data type.
VI-A5 Data Masking in Arrays
@@
We next evaluate the overhead of applying data access policies on array-typed columns where each element is of type struct. We consider two settings:
  • Unconditional masking: mask a field in each struct for all elements in the array.
  • Conditional masking: mask a field in each struct for elements in the array that satisfy a given condition. This results in an extra overhead of evaluating the condition for each element in the array.
Figure 10: Impact of array sizes on CPU overhead
@@
Figure 10 shows that the masking overhead grows linearly with array size for each field path. For a given array size, the overhead of unconditional masking is smaller than conditional masking due to the additional cost of comparing element values.
VI-B Data Guard in Production
Data Guard has been deployed as the default access control layer to serve data accesses to LinkedIn’s exabyte-scale data warehouse via Spark and Trino, ensuring compliance with our data usage policies and regulations. Table II shows some statistics on the usage of data-masking views in our production warehouses.
Based on case studies of some critical usecases, the average CPU overhead of data-masking views per application is 13.2%, which we consider acceptable given the importance of access control. Some data-masking views incur higher overhead when policies target multiple fields or deeply nested structures, consistent with our micro-benchmark results. By contrast, many views apply only row-level masking or mask a small number of top-level fields and therefore have much lower overhead. In some cases, row-level masking can even reduce query cost because filtering out rows decreases downstream work.
Metric Value
Number of purposes supported 50+
Number of unique policies enforced 110+
Number of tables protected with Data Guard policies 4,000+
Number of active data-masking views 13,000+
Number of protected daily table accesses 20,000+
Number of protected Spark applications 2,700+
Number of protected Trino applications 800+
Average number of policies applied per view 1.7
Maximum number of policies applied per view 27
Average depth of enforced fields 1.94
Maximum depth of enforced fields 9
Average CPU overhead of data-masking views per application 13.2%
TABLE II: Data Guard Production Statistics at LinkedIn
VII Practical Considerations
@@
In this section, we discuss several practical issues considered while deploying Data Guard at LinkedIn’s scale:
Enforcement Verification: Beyond access-time enforcement, Data Guard implements automated verification through continuous monitoring of storage access logs and query execution logs. This detects potential bypasses or accesses to stale views, providing comprehensive audit trails. Our automated remediation system notifies consumers of compliance issues and suggests corrective actions, closing the enforcement loop.
Policy Label Management: At LinkedIn, policy labels are carefully curated and organized into a hierarchical structure referred to as ontology. An ontology-based organization not only classifies data in our ecosystem, but also captures relationships between these data classes. Ontology also supports advanced use cases at LinkedIn related to reasoning and inference on compliance metadata. We note however that the access control model proposed in this paper does not require an ontology-based organization of policy labels.
Policy Management: At LinkedIn, purposes are organized into a directed acyclic graph (DAG), where edges denote parent–child relationships. A child may have multiple parents and inherits all parent policies, with conflicts resolved in favor of the more restrictive policy i.e. $MASK$ policies taking precedence over $KEEP$ policies. This hierarchy with inheritance reduces policy duplication across related purposes. Policies are managed like source code: they are versioned, code-reviewed, and synced to an online Policy Catalog that offers APIs for querying by purpose or label. Policies in Data Guard are intentionally simple, enabling software engineers and product managers—typically feature owners—to author them without requiring policy expertise.
Developer Experience: Data Guard fundamentally transforms how product teams interact with data compliance requirements. By abstracting compliance logic into centralized policies, teams no longer need specialized privacy knowledge or maintenance overhead of custom application-level implementations. This shift reduced development time from weeks to 2-3 days, as engineers can now focus on product features rather than interpreting complex compliance obligations. Our enhanced data catalog provides interactive masking impact visualizations and clear documentation of view behaviors.
@@
Engine Optimizations: Implementing masking via UDFs has a limitation that it prevents query optimizations like predicate pushdown or nested column pruning. Our focus so far has been on the APIs for enforcing policies and ensuring a friction-free developer experience. While current performance overhead remains acceptable for production workloads with optimizations implemented so far, we recognize opportunities for further optimizations.
VIII Conclusions And Future Work
@@
We have presented Data Guard - a fine-grained access control system for large data warehouses. Data Guard’s policy language supports specification of access-control rules based on semantic labels used to describe data present in the warehouse tables. Data Guard’s policy compiler translates these policies into SQL views which encode the data-masking logic. Further, Data Guard transparently routes table accesses to the appropriate data-masking views based on the accessor’s purpose, thus supporting purpose-based access control. Data Guard has been deployed at-scale in LinkedIn and is the primary policy enforcement mechanism for accesses to warehouse tables. We plan to continue work on a number of opportunities for performance optimizations in Data Guard as highlighted in the paper. We also plan to enhance Data Guard capabilities to support privacy-enhancing technologies such as anonymization and differential privacy, and encryption for data storage. Finally, defining a theoretical framework for the fine-grained access control problem and establishing the properties of Data Guard in the context of this framework is an important future direction.
IX Acknowledgment
@@
We would like to thank Kapil Surlaker and Raghu Hiremagalur for providing numerous critical inputs on the design of Data Guard. We would like to thank Yong Li, Manasa Subramanian, Divya Singh, Yanwen Lin, Justin Heaton, Vishwaa Patel, Bryan Ji, Qiang Fu, Steve Cao, Wenning Ding, and Carleen Li for their significant contributions to the development of Data Guard. Finally, we would like to thank Maneesh Varshney and Chris Harris for their helpful feedback in improving the quality of the paper.